# EF-3DGS: Event-Aided Free-Trajectory 3D Gaussian Splatting

### v3

@@ 4.3 Photometric Bundle Adjustment
$\mathcal{L}_{PBA}=\sum_{\mathbf{u}\in\mathcal{P}}\sum_{I_{s}\in\mathcal{F}}||I_{s}(\mathbf{u}^{\prime})-\hat{I}(\mathbf{u})||^{2},$ (15)
@@
Figure 3: The illustration of unified CMax and LEGM optimization. We warp previous event frames to the sampled timestamp through the optical flow and maximize the sharpness of the image of piece-wise warped events (IPWE). The byproduct IPWE is utilized to establish additional constraints on 3DGS.
@@
Table 1: Quantitative evaluations on Tanks and Temples. Each baseline method is trained with its public code under the original settings and evaluated with the same evaluation protocol. The best results are highlighted in bold.
@@
4.4 Fixed-GS Training Strategy
[…]
4.5 Overall Training Pipeline
Assembling all loss terms, we get the overall loss function:
where $\lambda_{PBA}$ are the weighting factor. Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We incorporate dynamic scene allocation strategies from LocalRF [28] for handling extended video sequences. Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS [9] while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction. Please refer to the Supplementary Material for the algorithm pipeline and additional implementation details.
5 Experiments
@@
Figure 5: Pose estimation comparison. We visualise the trajectory (3D plot) and $\textrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\textrm{RPE}_{r}$ by a quarter of the max $\textrm{RPE}_{r}$ across all results of each scene.
5.1 Dataset
Tanks and Temples.  We conduct comprehensive experiments on the Tanks and Temples dataset [22]. Similar to LocalRF [28], we adopt 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF [28], we apply 4× spatial downsampling to the videos. To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS. The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios. To synthesize realistic event data, we first upsample the original videos by [18] and then apply the simulator V2E [14].
RealEv-DAVIS.  Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS, comprising four outdoor scenes. Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS. We employ COLMAP for ground-truth poses. For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames. Further details are provided in the supplementary material.
5.2 Implementation details
We follow the optimization parameters by the configuration outlined in the 3DGS [20]. We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decays to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for the following CMax warping. The intervals of neighboring warping $r$ in CMax are set to 3. The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS. We provide detailed ablation studies on these hyperparameters and additional implementation details in the supplementary material.
@@ 5.3 Experimental Setup
Metrics.  We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM [49], and LPIPS [54]. For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics [55, 39], as delineated in [4]. Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
Baselines.  For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF [28] and F2-NeRF [48]. We also include pose-free methods like Nope-NeRF [4] and CF-3DGS [9]. (2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF [21], EvDeblurNeRF [5] and Event-3DGS [16]. Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens [43]) to temporally upsample frames before feeding them into CF-3DGS.
5.4 Experimental Results
We select every ten frames as a test image for NVS evaluation following LocalRF [28]. Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in iNeRF [53], we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
@@ 5.4 Experimental Results
Performance under Varying Camera Speeds  As shown in Fig 2 and Fig. 3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience. The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS [9] increasing from 0.61dB at 6FPS to 3.43dB at 1FPS. Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS). This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
@@ 5.5 Ablation Studies
Effect of Each Component  Table 4 presents a comprehensive ablation study of our key components under the challenging 1FPS setting on Tanks and Temples. $\mathcal{L}_{EGM}$ serves as the foundation of our approach, providing substantial improvements in both rendering quality (+1.63dB PSNR) and pose accuracy by enabling rich supervision between discrete frames. Building upon this, $\mathcal{L}_{LEGM}$ extracts motion information from events and constrains 3DGS in the gradient domain, significantly improving pose estimation while modestly enhancing rendering quality. $\mathcal{L}_{PBA}$, though designed to address color inconsistency issues, not only improves rendering quality but also enhances pose estimation accuracy by establishing geometric and photometric consistency across views. The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization. We provide more intuitive ablation visualizations in the supplementary material.
Figure 6: Robustness of different methods to pose disturbance.
Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF [23], we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 6. Notably, Event-3DGS [16], which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of joint pose-scene optimization. Furthermore, Our proposed framework demonstrates superior tolerance across all perturbation levels. Even under significant noise, our method experiences substantially less degradation in both rendering quality and trajectory accuracy.

### v4

@@ 4.3 Photometric Bundle Adjustment
[…]
@@
4.4 Fixed-GS Training Strategy
$G_{\theta}^{*}=\underset{f}{\operatorname{argmin}}~\mathcal{L}_{color},~t\in\{t_{i,0}\}$ (17)
4.5 Overall Training Pipeline
Assembling all loss terms, we get the overall loss function:
where $\lambda_{PBA}$ are the weighting factor. Note that since event cameras typically record only the changes in brightness intensity, the $\mathcal{L}_{EGM}$ and $\mathcal{L}_{LEGM}$ losses are computed in the grayscale domain, whereas the $\mathcal{L}_{PBA}$ loss is calculated in RGB color space. We incorporate dynamic scene allocation strategies from LocalRF localrf for handling extended video sequences. Our overall training pipeline builds upon the progressive optimization scheme of CF-3DGS cf3dgs while introducing novel components to integrate event stream data for robust free-trajectory scene reconstruction. Please refer to Section˜A.3 for the algorithm pipeline and additional implementation details.
5 Experiments
@@
Table 1: Quantitative evaluations on Tanks and Temples dataset. The best results are highlighted in bold. Note that, in the “input” column, “F” denotes traditional frame input, while “E+F” denotes hybrid frame and event input.
@@
5.1 Dataset
Tanks and Temples.  We conduct comprehensive experiments on the Tanks and Temples dataset tnt . Similar to LocalRF localrf , we adopt 9 scenes, covering large-scale indoor and outdoor scenes. For each scene, we sample a video clip with a 50-second duration, typically featuring free camera trajectories and covering a considerable distance. Following LocalRF localrf , we apply 4× spatial downsampling to the videos. To evaluate the robustness under varying camera speeds, we employ varying temporal downsampling of 6 FPS, 4 FPS, 3 FPS, 2 FPS, and 1 FPS. The reduction in frame rate effectively creates larger inter-frame displacements, simulating high-speed scenarios. To synthesize realistic event data, we first upsample the original videos by RIFE and then apply the simulator V2E v2e .
RealEv-DAVIS.  Due to the lack of free-trajectory event camera datasets, we introduce RealEv-DAVIS. Using a DAVIS346 camera that simultaneously captures frames and events at 346×260 resolution, we record 40-second handheld sequences at 25 FPS. We employ COLMAP for ground-truth poses. For SLOW scenarios, we retain every second frame, while for FAST scenarios, we keep only one frame per five frames. Further details are provided in Section˜A.2.
5.2 Implementation details
We follow the optimization parameters by the configuration outlined in the 3DGS 3dgs . We optimize the camera poses in the representation of quaternion rotation. The initial learning rate is set to $10^{-5}$ and gradually decays to $10^{-6}$ until convergence. The balancing weight $\lambda_{cm}$, $\lambda_{grad}$ and $\lambda_{PBA}$ is empirically set to 0.1, 0.2 and 0.5. For the division of events between adjacent frames, we maintain a constant interval of $\frac{1}{6}$s for Tanks and Temples and $\frac{1}{25}$s for RealEv-DAVIS, setting the number of subinterval N accordingly. For example, in Tanks and Temples, N equals 2 for 3FPS and 6 for 1FPS. This ensures adherence to the constant brightness assumption within each sub-interval and provides adequate events for the following CMax warping. The intervals of neighboring warping $r$ in CMax are set to 3. The contrast threshold $C$ is set to 0.25 for Tanks and Temples and 0.21 for RealEv-DAVIS. We provide detailed ablation studies on these hyperparameters and additional implementation details in Section˜A.4.
@@ 5.2 Implementation details
Figure 5: Pose estimation comparison. We visualise the trajectory (3D plot) and $\mathrm{RPE}_{r}$ (color bar) of each method. We clip and normalize the $\mathrm{RPE}_{r}$ by a quarter of the max $\mathrm{RPE}_{r}$ across all results of each scene.
@@ 5.3 Experimental Setup
Metrics.  We evaluate all the methods from two aspects: novel view synthesis and pose estimation. For the novel view synthesis task, we report the standard metrics PSNR, SSIM ssim , and LPIPS lpips . For the pose estimation task, we adopt the Absolute Trajectory Error (ATE) and Relative Pose Error (RPE) metrics pose_eval1 ; pose_eval2 , as delineated in nopenerf . Since these metrics are inherently influenced by frame rate, we upsample all estimated poses to a consistent temporal resolution before evaluation for fair comparison across different frame rate settings.
Baselines.  For a fair comparison, we focus on two categories of methods: (1) For frame-based approaches, we selected methods specifically addressing free-trajectory scenarios, such as LocalRF localrf and F2-NeRF f2nerf . We also include pose-free methods like Nope-NeRF nopenerf and CF-3DGS cf3dgs . (2) For event-frame hybrid methods, we consider approaches that fuse events and frames, including ENeRF enerf , EvDeblurNeRF evdeblurnerf and Event-3DGS event3dgs .
@@
Since no existing method integrates events for free-trajectory scenarios, we implement EvCF-3DGS as a competitive baseline that leverages an event-based frame interpolation network (Time Lens timelens ) to temporally upsample frames before feeding them into CF-3DGS.
5.4 Experimental Results
We select every ten frames as a test image for NVS evaluation following LocalRF localrf . Since the camera poses are unknown in our setting, we need to estimate the poses of test views. As in iNeRF iNeRF , we freeze the 3DGS model, initialize the test poses with the poses of the nearest training frames, and optimize the test poses by minimizing the photometric error between rendered images and test views.
@@ 5.4 Experimental Results
Figure 6: Robustness of different methods to pose disturbance.
Performance under Varying Camera Speeds  As shown in Table 2 and Table 3, while all methods degrade as the frame rate decreases, our approach shows remarkable resilience. The performance gap widens significantly at lower frame rates, with our PSNR advantage over CF-3DGS cf3dgs increasing from 0.61dB at 6FPS to 3.43dB at 1FPS. Notably, our method also consistently outperforms other event-based methods (EvCF-3DGS and Event-3DGS). This confirms not only the value of event data in challenging scenarios but also the superiority of our integration approach.
@@ 5.5 Ablation Studies
Effect of Each Component  Table 4 presents a comprehensive ablation study of our key components under the challenging 1FPS setting on Tanks and Temples. $\mathcal{L}_{EGM}$ serves as the foundation of our approach, providing substantial improvements in both rendering quality (+1.63dB PSNR) and pose accuracy by enabling rich supervision between discrete frames. Building upon this, $\mathcal{L}_{LEGM}$ extracts motion information from events and constrains 3DGS in the gradient domain, significantly improving pose estimation while modestly enhancing rendering quality. $\mathcal{L}_{PBA}$, though designed to address color inconsistency issues, not only improves rendering quality but also enhances pose estimation accuracy by establishing geometric and photometric consistency across views. The Fixed-GS training strategy, while having no impact on pose optimization, significantly improves rendering quality by effectively separating structure and color optimization. We provide more intuitive ablation visualizations in Section˜A.5.
Robustness to Pose Disturbance  To validate the robustness of different methods under inaccurate pose initialization, a common challenge in practical scenarios, we introduce varying degrees of perturbations to the initial camera poses estimated by COLMAP. Specifically, following BARF barf , we parametrize the camera poses $\mathbf{p}$ with the $\mathfrak{se}(3)$ Lie algebra. For each scene, we synthetically perturb the camera poses with additive noise $\delta\mathbf{p}\sim\mathcal{N}(\mathbf{0},n\mathbf{\textrm{I}})$, where n is the noise level. Then, each method is initialized with the noised poses, after which the optimization is performed. The results are illustrated in Fig. 6. Notably, Event-3DGS event3dgs , which lacks the capability to optimize camera poses, exhibits a drastic performance degradation as the magnitude of pose disturbances increases. This observation validates the critical importance of joint pose-scene optimization. Furthermore, Our proposed framework demonstrates superior tolerance across all perturbation levels. Even under significant noise, our method experiences substantially less degradation in both rendering quality and trajectory accuracy.